OpenStudy (tia_sweetheart17):

solve for x: 2/3x=10

OpenStudy (anonymous):

you multiply bu 3/2 on both sides

OpenStudy (shamil98):

\[\large \frac{ 2 }{ 3 }x = 10\] \[\large \frac{ 3 }{ 2 } \times \frac{ 2 }{ 3 } x = 10 \times \frac{ 3 }{ 2 }\] \[\huge x = 5\]